OpenStudy (anonymous):

f(t) ={ (0, t<4) , ( (t^2 -8t + 20), t=>4) } find the laplace transform

OpenStudy (anonymous):

the unit step function must be constructed from \[0, t<4\] and \[t^2 - 8t + 20, t \ge4\]

OpenStudy (anonymous):

it would start at 0 +u(t-4)(t^2 -8t + 20) right?

OpenStudy (ash2326):

Yeah f(t) can be written as \[f(t)= (t^2-8t+20) U(t-4)\]

OpenStudy (ash2326):

You need to find the LT???

OpenStudy (anonymous):

right... does the laplace transform just distribute over into each term?

OpenStudy (ash2326):

Yeah it can be distributed but you know that LT is defined as \[f(s)=\int_{0}^{\infty} f(t) e^{-st} dt\] so we'll have to use a LT propertu

OpenStudy (ash2326):

*property

OpenStudy (ash2326):

You know that property?

OpenStudy (anonymous):

err no..

OpenStudy (ash2326):

Ok \[\large LT( f(t) U(t-a))= e^{-as} F(s)\] Have you studied this? Do you want me to derive this?

OpenStudy (anonymous):

ooh yes i know that

OpenStudy (ash2326):

Ok so we'll use it First we'll find the LT of \[t^2-8t+20\] Can you find that?

OpenStudy (anonymous):

2/(s^3) + -8/(s^2) + 20/s

OpenStudy (anonymous):

from the table

OpenStudy (ash2326):

Great:) Now just we need to multiply it by \[\large e^{-4s}\] and you have your LT \[\huge e^{-4s} ( \frac{2}{s^3}-\frac{8}{s^2}+\frac{20 }{s})\]

OpenStudy (anonymous):

but the answer contains terms (1+4s)8 for the middle term, and (1+4s+8s^2) on the first in the answer book

OpenStudy (ash2326):

It can also be written as \[\huge e^{-4s} \frac{2-8s+20s^2}{s^3}\] But it's also not not matching!!! Could you post the complete answer

OpenStudy (anonymous):

sure one sec

OpenStudy (ash2326):

Oh no I made a mistake

OpenStudy (ash2326):

Don't post the answer, we'll get it

OpenStudy (ash2326):

Actually the property is \[\huge LT( f(t-a) u(t-a))=e^{-as} F(s)\]

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

sidetrack: this would work with sin(pit) right? (my next question)

OpenStudy (ash2326):

We have \[f(t)= t^2-8t-20\] so \[f(t-4)=t^2-8t+16-8t+32-20=t^2-16t+28\] so f(t) can be written as \[f(t)= f(t-4)+8t-48\] Do you understand this?

OpenStudy (anonymous):

oh you plugged in t-4 into all the t's

OpenStudy (ash2326):

yeah

OpenStudy (ash2326):

You understand the last step, where I write f(t) in terms of f(t-4)??

OpenStudy (anonymous):

on second thought no...

OpenStudy (anonymous):

OH WAIT

OpenStudy (anonymous):

you added and subtracted 48 to make it equal to our original function

OpenStudy (anonymous):

added 8t rather

OpenStudy (ash2326):

Yeah but made a mistake

OpenStudy (ash2326):

f(t)=t^2-8t+20 so \[f(t-4)=t^2+16-8t-8t+32+20=t^2-16t+68\] so \[f(t)=f(t-4)+8t-48\]

OpenStudy (anonymous):

righto

OpenStudy (ash2326):

Great:D So I can further change this as \[\huge f(t)=f(t-4)+8(t-4)-16\] You got this???

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

8(t-4) -16 = 8t -48 ?

OpenStudy (ash2326):

Yeah !!!! Okay so \[LT(f(t)U(t-4))=LT(f(t-4)+8(t-4)-16)U(t-4))\] \[= e^{-4s}(LT(f(t))+8 LT(t)-16 LT(1)) \]

OpenStudy (anonymous):

what is the point of doing that?

OpenStudy (ash2326):

Because we want to use \[LT[f(t-4) U(t-4)]= e^{-4s} F(s)\]

OpenStudy (anonymous):

ah so t-4 will go into each t ok

OpenStudy (ash2326):

Yeah!!!

OpenStudy (ash2326):

Now you find LT[f(t)] f(t)= t^2-8t+20 and substitute in above expression

OpenStudy (anonymous):

so L(f(t-4)) + L(8(t-4)) so on so forth?

OpenStudy (ash2326):

Nope we have \[e^{-4s} ( LT (f(t)+8LT(t)-16 LT(1))\]

OpenStudy (ash2326):

We have used the property which I mentioned earlier

OpenStudy (ash2326):

Did you understand?

OpenStudy (anonymous):

kinda im at this point \[e ^{-4s} (L(ft))\]

OpenStudy (anonymous):

f(t) being f(t-4) +8(t-4) -16

OpenStudy (anonymous):

ugh im lost

OpenStudy (ash2326):

No here for LT of f(t-4) use LT of f(t) for LT of (t-4) use LT( t) and -16 is just -16/s

OpenStudy (amistre64):

just curious if this works: since 0 to 4 is zero we can ignore that part \[ \int_{4}^{inf} e^{-st}(t^2 -8t + 20) dt=...\] e^-st t^2 -8t +20 e^-st/-s -(2t -8) e^-st/s^2 2 e^-st/-s^3 at inf we go zero sooo, if i push my negative thru t^2 -8t +20 e^-st/s 2t -8 e^-st/s^2 2 e^-st/s^3 16 -32 +20 e^-4s/s 8 -8 e^-4s/s^2 2 e^-4s/s^3 e^-4s (4/s+ 2/s^3) But then I aint got a clue as to what your doing with that U stuff

OpenStudy (ash2326):

@amistre64 This would work always I'm trying to create the form \[f(t-4) U(t-4)\] where U is the unit step function so that I could use the property and write Laplace transform as \[e^{-4s}F(s)\]

OpenStudy (amistre64):

:) yeah, I got no idea how to do it like that

OpenStudy (ash2326):

I can write f(t) as \[f(t)U(t-4)= (f(t-4)+8(t-4)-16) U(t-4)\] Taking LT now \[e^{-4s} (F(s)+ 8/s^2-16/s)\] \[F(s)=LT(f(t)= 2/s^3-8/s^2+20/s\] so we get \[e^{-4s}(2/s^3+4/s)\] I'm making some mistake, I'll check again

OpenStudy (ash2326):

oh wait you also got that!!!

OpenStudy (amistre64):

i did :) but i make no gaurentees for its correctness ;)

OpenStudy (ash2326):

Oh can you bump this!!!

OpenStudy (amistre64):

it says bump in 10min

OpenStudy (ash2326):

But it's come to the top. Thanks:)

OpenStudy (amistre64):

oh, then that means it can bump again in 10 mins :)

OpenStudy (ash2326):

Yeah:)

OpenStudy (phi):

Here's two ways